OpenStudy (anonymous):

any body help me how to solve lim x->0 (sin4x)/x

hartnn (hartnn):

follow your last post.

OpenStudy (anonymous):

i cant understand how>?

hartnn (hartnn):

do u know this standard limit : \(\lim \limits_{x \rightarrow0}\frac{\sin x}{x}=1\)

OpenStudy (anonymous):

yes i konw that

OpenStudy (anonymous):

i cant use l hospital rules

OpenStudy (anonymous):

in my book answer is 4 i dont know how??

hartnn (hartnn):

so, do this , multiply the numerator and denominator by 4 \(\huge 4\lim \limits_{x \rightarrow0}\frac{\sin 4x}{4x}=...?\)

hartnn (hartnn):

\(\huge 4\lim \limits_{4x \rightarrow0}\frac{\sin 4x}{4x}=...?\) now put 4x = y and you get a standard limit whose value is 1 (with 4)

OpenStudy (anonymous):

if i multiply that what did i got?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin\] i got

hartnn (hartnn):

what! ? how ??

OpenStudy (anonymous):

can you give me full steps:

hartnn (hartnn):

let y=4x \(\large 4\lim \limits_{4x \rightarrow0}\frac{\sin 4x}{4x}=4[\lim \limits_{y \rightarrow0}\frac{\sin y}{y}]= 4(..?..)\)

OpenStudy (anonymous):

1you mean?

hartnn (hartnn):

yes, ?=1 so, 4*1=4

OpenStudy (anonymous):

\[4\lim4x→0\sin4x4x=4[limy→0sinyy]=4(1)= than is equal \to 4\]

OpenStudy (anonymous):

thank you soo much

hartnn (hartnn):

hmm...welcome ^_^